OpenStudy (anonymous):

Let f(x)=x^2-x-9 and g(x)=-4x-7. Find g(2) and f(g(2))

OpenStudy (anonymous):

g(2) = -4(2) - 7 = -15 f(g(2)) = (-15)^2 - (-15) - 9 = 231

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

u r welcome...